# Robust Tail Index Estimation under Random Censoring via Minimum Density Power Divergence

### v5

@@ Abstract
We propose a robust estimator for the tail index of Pareto-type distributions under random right-censoring, based on the minimum density power divergence (MDPD) framework. To our knowledge, this is the first application of the MDPD approach to extreme value models with random censoring, opening a new direction for robust inference in this setting. Under mild regularity conditions, the estimator is shown to be consistent and asymptotically normal. Its performance in finite samples is extensively evaluated through simulation studies, demonstrating superior robustness and efficiency compared to existing methods. Contamination is introduced only before censoring to provide a meaningful assessment of robustness, while contamination after censoring is shown to yield distorted or unrealistic results. The practical relevance of the approach is illustrated using a real dataset on insurance claims, which features light censoring and fully observable extremes, and a dataset on AIDS survival times, which, despite stronger censoring $(p<1/2),$ allows for illustrative comparisons and highlights practical limitations and challenges in more difficult scenarios.
@@ 1. Introduction
which determines whether or not $X$ has been observed. We assume that the tail functions $\overline{F}:=1-F$ and $\overline{G}:=1-G$ are regularly varying at infinity (or Pareto-type) of positive tail indices $\gamma_{1}>0$ and $\gamma_{2}>0,$ allowing for slowly varying deviations from the strict Pareto form. More precisely, for any $x>0,$
@@
$\widehat{\gamma}_{1}^{\left(EFG\right)}:=\frac{\widehat{\gamma}^{\left(H\right)}}{\widehat{p}}$,
@@
$\widehat{\gamma}^{\left(H\right)}:=k^{-1}{\displaystyle\sum\limits_{i=1}^{k}}\log\left(Z_{n-i+1:n}/Z_{n-k:n}\right),$
@@
The integer sequence $k=k_{n}$ represents the number of top order statistics used in the estimation of the tail index so that $k\rightarrow\infty$ and $k/n\rightarrow 0$ as $n\rightarrow\infty.$ This ensures that enough extreme observations are included for reliable estimation, while avoiding the inclusion of too many moderate values that could bias the tail index estimate.
@@
$\widehat{\gamma}_{1}^{\left(W\right)}:={\displaystyle\sum\limits_{i=1}^{k}}\frac{\overline{F}_{n}^{\left(KM\right)}\left(Z_{n-i:n}\right)}{\overline{F}_{n}^{\left(KM\right)}\left(Z_{n-k:n}\right)}\log\frac{Z_{n-i+1:n}}{Z_{n-i:n}},$
@@
$F_{n}^{\left(KM\right)}\left(x\right):=\underset{Z_{i:n}\leq x}{1-{\displaystyle\prod}}\left(\dfrac{n-i}{n-i+1}\right)^{\delta_{\left[i:n\right]}},$
denotes the popular Kaplan-Meier estimator of cdf $F$ (Kaplan and Meier, 1958). The asymptotic normality of $\widehat{\gamma}_{1}^{\left(W\right)}$ is established in Beirlant et al. (2019) by considering Hall’s model (Hall, 1982). Bias reduction in tail index estimation under censoring has been addressed in Beirlant et al. (2016), Beirlant et al. (2018) and Beirlant et al. (2019).
By using a Nelson-Aalen integration, recently Meraghni et al. (2025) derived a new estimator for $\gamma_{1}$ given by
$\widehat{\gamma}_{1}^{\left(MNS\right)}:={\displaystyle\sum\limits_{i=1}^{k}}\frac{\delta_{\left[n-i+1:n\right]}}{i}\frac{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-i+1:n}\right)}{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-k:n}\right)}\log\frac{Z_{n-i+1:n}}{Z_{n-k:n}},$
@@
$F_{n}^{\left(NA\right)}\left(x\right)=1-\prod_{Z_{i:n}<x}\exp\left\{-\frac{\delta_{\left[i:n\right]}}{n-i+1}\right\},$ (1.4)
denotes the well-known Nelson-Aalen estimator of cdf $F$ (Nelson, 1972). Note that
$\frac{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-i+1:n}\right)}{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-k:n}\right)}=\prod_{j=i+1}^{k}\exp\left\{-\frac{\delta_{\left[n-j+1:n\right]}}{j}\right\},$ (1.5)
thus the formula of $\widehat{\gamma}_{1}^{\left(MNS\right)}$ can be rewritten as
$\widehat{\gamma}_{1}^{\left(MNS\right)}:={\displaystyle\sum\limits_{i=1}^{k}}a_{ik}\log\frac{Z_{n-i+1:n}}{Z_{n-k:n}},$
@@
The asymptotic properties of both the Kaplan-Meier-based estimator $\widehat{\gamma}_{1}^{\left(W\right)}$ and the Nelson-Aalen-based estimator $\widehat{\gamma}_{1}^{\left(MNS\right)}$ are established provided that $p>1/2.$ This condition ensures that a sufficiently large fraction of extreme observations remains uncensored, which is essential for reliable estimation of $\gamma_{1}.$ Equivalently, it can be expressed as $\gamma_{1}<\gamma_{2},$ meaning that the censoring distribution has a lighter tail than the distribution of interest. If $p\leq 1/2$ (or $\gamma_{1}\geq\gamma_{2}$), the effective tail sample size may be too small, and the Gaussian approximation underlying these estimators may break down. In practice, this condition serves as a natural guideline for the applicability of both estimators, emphasizing the importance of the relative tail heaviness between the observed and censoring distributions.
@@
The most common method for estimating the parameters of an extreme value distribution in extreme value analysis relies on the maximum likelihood estimation (MLE) method. Although these estimators possess desirable asymptotic properties, they can be sensitive to outlying observations from the assumed extreme value models; see, for instance, Brazauskas and Serfling (2000). Consequently, robust statistical methods provide a better alternative to mitigate the influence of outliers and deviations from the underlying parametric models. It has been shown that employing robust statistical ideas in extreme value theory improves the quality and precision of estimates (Dell’Aquila and Embrechts, 2006).
Juárèz and Schucany (2004) appear to be the first authors to employ the minimum density power divergence (MDPD) of Basu et al. (1998) for the robust estimation of the parameters of an extreme value distribution. Since then, this divergence measure has become a widely used tool for robust estimation of parameters of extreme value distributions. Kim and Lee (2008), Dierchx et al. (2013), Goegebeur et al. (2014), Dierckx et al. (2021) have applied the MDPD approach to estimate the tail index and quantiles from Pareto-type distributions. Recently, Ghosh (2017) proposed a robust MDPD estimator for a real-valued tail index. This estimator is a robust generalization of the estimator proposed by Mattys and Beirlant (2003) and addresses non-identical distributions in the exponential regression model using the approach in Ghosh and Basu (2013). Furthermore, Dierchx et al. (2013) employed the MDPD concept on an extended Pareto distribution for relative excesses over a high threshold. Minkah el al. (2023A) and Minkah el al. (2023B) recently proposed a robust estimator for the tail index of a Pareto-type distribution using the MDPD approach applied to an exponential regression model.
@@
To the best of our knowledge, apart from the recent work of Dierckx et al. (2021), there exists no robust estimator specifically tailored for the tail index under random censoring. Their approach focuses on the conditional tail index, employing kernel smoothing techniques and local threshold selection for each covariate value. While the estimator effectively reduces bias and variance in the presence of contamination, it relies on a separate selection of tuning parameters and is restricted to the conditional setting, which complicates its direct application to the unconditional tail index. This limitation underscores the need for a robust and flexible framework capable of handling both unconditional and censored extreme value scenarios, motivating the development of the MDPD-based estimator proposed in this work.
@@
Finally, Appendix A compiles several technical lemmas and propositions that are instrumental to the theoretical development, while Appendix B contains the figures and additional results related to the simulation study.
@@ 2. Minimum density power divergence
Basu et al. (1998) introduced a new measure of divergence between two probability densities $\ell$ and $f,$ known as the density power divergence (DPD), by
@@ 2.1. MDPD estimator of $\gamma_{1}$ for right-censored data
$dF_{n}^{\left(NA\right)}\left(x\right)=\frac{\overline{F}_{n}^{\left(NA\right)}\left(x\right)dH_{n}^{\left(1\right)}\left(x\right)}{\overline{H}_{n}\left(x^{-}\right)},$ (2.11)
@@ 4. Simulation Study
In this section, we investigate the finite-sample performance of the proposed MDPD tail index estimator $\widehat{\gamma}_{1,\alpha}$ through an extensive Monte Carlo simulation study. The main objective is to assess both efficiency and robustness under random right censoring and contamination affecting the upper tail.
More specifically, the simulation design is intended to highlight the trade-off between efficiency and robustness induced by the tuning parameter $\alpha$ across a wide range of realistic and adverse configurations.
4.1. Data-generating mechanism
We consider two heavy-tailed distributions for the latent variable $X:$
  • Burr distribution: $F(x)=1-(1+x^{1/\eta})^{-\eta/\gamma_{1}},$ $x>0,$ with shape parameter $\eta=0.25$ and tail index $\gamma_{1}\in\{0.3,0.5\}.$
  • Fréchet distribution: $F(x)=\exp(-x^{-1/\gamma_{1}}),$ $x>0,$ with tail index $\gamma_{1}\in\{0.3,0.5\}.$
Right censoring is generated independently via a rv $C$ from the same parametric family as $F$ with censoring tail index $\gamma_{2},$ such that the proportion of uncensored extreme observations $p>1/2.$ This ensures that the observed rv $Z=\min(X,C)$ remains regularly varying and satisfies the weak censoring condition discussed in Section 1.
This construction guarantees a controlled censoring mechanism while preserving the asymptotic tail behavior of the observed data.
@@ 4.2. Contamination schemes
To evaluate robustness, three contamination scenarios are considered, reflecting increasing levels of difficulty and realism.
  1. (1)
    Original (uncontaminated) data:
    The latent rvs $X_{i}$ are generated from the target cdf $F,$ and the observed pairs $(Z_{i},\delta_{i})$ through the censoring mechanism. This setting provides a baseline for assessing efficiency under the correctly specified model.
  2. (2)
    Contamination before censoring:
    A fraction $\epsilon$ of the latent rvs $X_{i}$ is replaced by extreme values drawn from a Burr distribution $F_{c}$ with a heavier tail than the target cdf $F.$ The censoring mechanism is then applied to the contaminated latent sample, yielding the observed data
    $(Z_{i}^{(\epsilon)},\delta_{i}^{(\epsilon)}):=\left(\min(X_{i}^{(\epsilon)},C_{i}),\mathbb{I}_{\{X_{i}^{(\epsilon)}\leq C_{i}\}}\right),\text{ }i=1,\dots,n.$
    This scenario corresponds to a classical robustness framework in which contamination directly affects the true underlying distribution $F.$ The Burr distribution is chosen as a flexible heavy-tailed contaminating model, ensuring a pronounced impact on the upper tail while allowing precise control of tail heaviness. Although other heavy-tailed distributions could be considered, the Burr family provides a convenient parametrization and has been widely used in extreme value simulation studies.
    This setting serves as a benchmark for evaluating robustness when contamination acts at the latent level and is partially filtered by the censoring mechanism.
  3. (3)
    Contamination after censoring: only the observed pairs $(Z_{i},\delta_{i})$ are available. Pseudo-samples $(\widetilde{X}_{i})$ are first generated from the Kaplan–Meier estimator $F_{n}^{(KM)}$ using the inversion method, and contamination is applied to $(\widetilde{X}_{i}).$ The observed data are then reconstructed as
[…]
    This approach provides an illustrative robustness assessment in situations where contamination directly affects the observed extremes, rather than the latent variable itself. It reflects a realistic and particularly adverse scenario encountered in practice, where outliers may arise from data recording errors, reporting issues, or abnormal external events occurring after censoring has taken place.
4.3. Generation of pseudo-samples from the Kaplan–Meier estimator
When the latent variables $X_{i}$ are unobserved due to right censoring, pseudo-observations can be generated from the Kaplan–Meier estimator $F_{n}^{(KM)},$ constructed on the basis of the observed pairs $\left(Z_{i},\delta_{i}\right),$ using the inversion method.
Formally, let $U_{i},$ $i=1,\dots,n,$ be independent rvs uniformly distributed on $\left(0,1\right),$ defined on the probability space $\left(\Omega,\mathcal{A},\mathbf{P}\right).$ We define the pseudo-observations by
$\widetilde{X}_{i}=Q_{n}^{(KM)}(U_{i}),$
where
$Q_{n}^{(KM)}\left(s\right):=\min\left\{x\in\mathbb{R}:F_{n}^{(KM)}\left(x\right)\geq s\right\},\text{ }0<s<1,$
denotes the empirical quantile function associated with $F_{n}^{(KM)}.$
The resulting pseudo-sample can be interpreted as a sample drawn from an estimate of the latent distribution function $F,$ reconstructed from the censored data. This approach provides a principled way to approximate the underlying distribution while fully accounting for the censoring mechanism, and allows contamination to be introduced in a manner consistent with the information actually available to the statistician. As a consequence, the contamination mechanism remains coherent with the observed-data structure, which is essential for a meaningful robustness analysis.

### v6

@@ Abstract
We propose a robust estimator for the tail index of Pareto-type distributions under random right-censoring, constructed within the minimum density power divergence (MDPD) framework and based on the Nelson–Aalen estimator of the cumulative hazard function. To our knowledge, this is the first application of the MDPD methodology to tail index estimation in the presence of random censoring. Under mild regularity conditions and within the weak censoring regime, the estimator is shown to be consistent and asymptotically normal. Its finite-sample performance is assessed through Monte Carlo simulations, revealing improved robustness–efficiency trade-offs compared to standard non-robust tail index estimators. Robustness is investigated under both pre-censoring and post-censoring contamination schemes. While pre-censoring contamination provides a meaningful framework for robustness assessment, post-censoring contamination directly alters the observable data and highlights the sensitivity of reconstruction-based approaches. The practical relevance of the method is illustrated using an insurance claims dataset with light censoring and fully observable extremes. An additional application to AIDS survival data is included for illustrative purposes, emphasizing the challenges encountered under stronger censoring.
@@ 1. Introduction
which determines whether or not $X$ has been observed. We assume that the tail functions $\overline{F}:=1-F$ and $\overline{G}:=1-G$ are regularly varying at infinity (or Pareto-type) with positive tail indices $\gamma_{1}>0$ and $\gamma_{2}>0,$ allowing for slowly varying deviations from the strict Pareto form. More precisely, for any $x>0,$
@@
$\widehat{\gamma}_{1}^{\left(\mathrm{EFG}\right)}:=\frac{\widehat{\gamma}^{\left(H\right)}}{\widehat{p}}$,
@@
$\widehat{\gamma}^{\left(\mathrm{H}\right)}:=k^{-1}{\displaystyle\sum\limits_{i=1}^{k}}\log\left(Z_{n-i+1:n}/Z_{n-k:n}\right),$
@@
The integer sequence $k=k_{n}$ represents the number of top order statistics used in the estimation of the tail index, such that $k\rightarrow\infty$ and $k/n\rightarrow 0$ as $n\rightarrow\infty.$ This ensures that enough extreme observations are included for reliable estimation, while avoiding the inclusion of too many moderate values that could bias the tail index estimation.
@@
$\widehat{\gamma}_{1}^{\left(\mathrm{W}\right)}:={\displaystyle\sum\limits_{i=1}^{k}}\frac{\overline{F}_{n}^{\left(KM\right)}\left(Z_{n-i:n}\right)}{\overline{F}_{n}^{\left(KM\right)}\left(Z_{n-k:n}\right)}\log\frac{Z_{n-i+1:n}}{Z_{n-i:n}},$
@@
$F_{n}^{\left(\mathrm{KM}\right)}\left(x\right):=\underset{Z_{i:n}\leq x}{1-{\displaystyle\prod}}\left(\dfrac{n-i}{n-i+1}\right)^{\delta_{\left[i:n\right]}},$
denotes the popular Kaplan-Meier estimator of the cdf $F$ (Kaplan and Meier, 1958). The asymptotic normality of $\widehat{\gamma}_{1}^{\left(\mathrm{W}\right)}$ is established in Beirlant et al. (2019) by considering Hall’s model (Hall, 1982). Bias reduction in tail index estimation under censoring has been addressed in Beirlant et al. (2016), Beirlant et al. (2018) and Beirlant et al. (2019).
By using a Nelson-Aalen integration, more recently Meraghni et al. (2025) derived a new estimator for $\gamma_{1}$ given by
$\widehat{\gamma}_{1}^{\left(\mathrm{MNS}\right)}:={\displaystyle\sum\limits_{i=1}^{k}}\frac{\delta_{\left[n-i+1:n\right]}}{i}\frac{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-i+1:n}\right)}{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-k:n}\right)}\log\frac{Z_{n-i+1:n}}{Z_{n-k:n}},$
@@
$F_{n}^{\left(\mathrm{NA}\right)}\left(x\right)=1-\prod_{Z_{i:n}<x}\exp\left\{-\frac{\delta_{\left[i:n\right]}}{n-i+1}\right\},$ (1.4)
denotes the well-known Nelson-Aalen estimator of the cdf $F$ (Nelson, 1972). Note that
$\frac{\overline{F}_{n}^{\left(\mathrm{NA}\right)}\left(Z_{n-i+1:n}\right)}{\overline{F}_{n}^{\left(\mathrm{NA}\right)}\left(Z_{n-k:n}\right)}=\prod_{j=i+1}^{k}\exp\left\{-\frac{\delta_{\left[n-j+1:n\right]}}{j}\right\},$ (1.5)
thus the formula of $\widehat{\gamma}_{1}^{\left(\mathrm{MNS}\right)}$ can be rewritten as
$\widehat{\gamma}_{1}^{\left(\mathrm{MNS}\right)}:={\displaystyle\sum\limits_{i=1}^{k}}a_{ik}\log\frac{Z_{n-i+1:n}}{Z_{n-k:n}},$
@@
The asymptotic properties of both the Kaplan-Meier-based estimator $\widehat{\gamma}_{1}^{\left(\mathrm{W}\right)}$ and the Nelson-Aalen-based estimator $\widehat{\gamma}_{1}^{\left(\mathrm{MNS}\right)}$ are established provided that $p>1/2.$ This condition ensures that a sufficiently large fraction of extreme observations remains uncensored, which is essential for reliable estimation of $\gamma_{1}.$ Equivalently, it can be expressed as $\gamma_{1}<\gamma_{2},$ meaning that the censoring distribution has a lighter tail than the distribution of interest. If $p\leq 1/2$ (or $\gamma_{1}\geq\gamma_{2}$), the effective tail sample size may be too small, and the Gaussian approximation underlying these estimators may break down. In practice, this condition serves as a natural guideline for the applicability of both estimators, emphasizing the importance of the relative tail heaviness between the observed and censoring distributions.
@@
The most common method for estimating the parameters of an extreme value distribution in extreme value analysis relies on maximum likelihood estimation (MLE) method. Although these estimators possess desirable asymptotic properties, they can be sensitive to outlying observations from the assumed extreme value models; see, for instance, Brazauskas and Serfling (2000). Consequently, robust statistical methods provide a better alternative to mitigate the influence of outliers and deviations from the underlying parametric models. It has been shown that employing robust statistical ideas in extreme value theory improves the quality and precision of estimates (Dell’Aquila and Embrechts, 2006).
Juárèz and Schucany (2004) appear to be the first authors to employ the minimum density power divergence (MDPD) of Basu et al. (1998) for the robust estimation of the parameters of an extreme value distribution. Since then, this divergence measure has become a widely used tool for robust estimation of parameters of extreme value distributions. Kim and Lee (2008), Dierchx et al. (2013), Goegebeur et al. (2014), and Dierckx et al. (2021) have applied the MDPD approach to estimate the tail index and quantiles from Pareto-type distributions. Recently, Ghosh (2017) proposed a robust MDPD estimator for a real-valued tail index. This estimator is a robust generalization of the estimator proposed by Mattys and Beirlant (2003) and addresses non-identical distributions in the exponential regression model using the approach in Ghosh and Basu (2013). Furthermore, Dierchx et al. (2013) employed the MDPD concept on an extended Pareto distribution for relative excesses over a high threshold. Minkah el al. (2023A) and Minkah el al. (2023B) recently proposed a robust estimator for the tail index of a Pareto-type distribution using the MDPD approach applied to an exponential regression model.
@@
To the best of our knowledge, apart from the recent work of Dierckx et al. (2021), there exists no robust estimator specifically tailored to the tail index under random censoring. Their approach focuses on the conditional tail index, employing kernel smoothing techniques and local threshold selection for each covariate value. While the estimator effectively reduces bias and variance in the presence of contamination, it relies on a separate selection of tuning parameters and is restricted to the conditional setting, which complicates its direct application to the unconditional tail index. This limitation underscores the need for a robust and flexible framework capable of handling both unconditional and censored extreme value scenarios, motivating the development of the MDPD-based estimator proposed in this work.
@@
Finally, Appendix A compiles several technical lemmas and propositions that are instrumental for the theoretical development, while Appendix B contains the figures and additional results related to the simulation study.
@@ 2. Minimum density power divergence
Basu et al. (1998) introduced a new measure of divergence between two probability densities $\ell$ and $f,$ known as the density power divergence (DPD), defined by
@@ 2.1. MDPD estimator of $\gamma_{1}$ for right-censored data
$dF_{n}^{\left(\mathrm{NA}\right)}\left(x\right)=\frac{\overline{F}_{n}^{\left(\mathrm{NA}\right)}\left(x\right)dH_{n}^{\left(1\right)}\left(x\right)}{\overline{H}_{n}\left(x^{-}\right)},$ (2.11)
@@ 4. Simulation Study
In this section, we investigate the finite-sample performance of the proposed MDPD tail index estimator $\widehat{\gamma}_{1,\alpha}$ through an extensive Monte Carlo simulation study. The main objective is to assess both efficiency and robustness in the presence of random right censoring and contamination affecting the upper tail of the distribution. Particular attention is paid to the impact of contamination on bias, MSE, and stability with respect to the choice of the number of upper order statistics. The simulation design is chosen so as to remain fully consistent with the theoretical framework developed in the previous sections.
4.1. Compared estimators and efficiency–robustness trade-off
The MDPD estimator $\widehat{\gamma}_{1,\alpha}$ is computed for $\alpha\in\{0.1,0.3,0.5\},$ allowing us to explore the efficiency–robustness trade-off induced by the tuning parameter. Small values of $\alpha$ emphasize efficiency under correct model specification, whereas larger values enhance robustness against contamination in the upper tail.
For comparison purposes, the following benchmark estimators are included:
  • the adapted Hill-type estimator $\widehat{\gamma}_{1}^{(\mathrm{EFG})}$,
  • the Kaplan–Meier-based estimator $\widehat{\gamma}_{1}^{(\mathrm{W})},$
  • the Nelson–Aalen-based estimator $\widehat{\gamma}_{1}^{(\mathrm{MNS})},$ corresponding to the special case $\widehat{\gamma}_{1,0}.$
This selection encompasses both classical and censoring-adapted tail index estimators that are standard in the extreme value literature. It therefore provides a relevant and informative benchmark for assessing the efficiency and robustness of the proposed MDPD-based estimator under random censoring.
4.2. Simulation models and censoring design
We consider two classical heavy-tailed distributions for the latent lifetime variable $X,$ which are widely used as benchmark models in extreme value analysis.
The first model is the Burr distribution, defined by
$F(x)=1-\left(1+x^{1/\eta}\right)^{-\eta/\gamma_{1}},\qquad x>0,$
with shape parameter $\eta=0.25$ and tail index $\gamma_{1}\in\{0.3,0.5\}.$
The second model is the Fréchet distribution, defined by
$F(x)=\exp\left(-x^{-1/\gamma_{1}}\right),\qquad x>0,$
with tail index $\gamma_{1}\in\{0.3,0.5\}.$
Right censoring is generated independently through a random variable $C$ drawn from the same parametric family as $X,$ with censoring tail index $\gamma_{2}.$ The model parameters are chosen such that the proportion of uncensored extreme observations satisfies $p>1/2.$ This condition ensures that the observed variable $Z=\min(X,C)$ remains regularly varying and fulfills the weak censoring condition introduced in Section 1. This construction allows us to control the censoring mechanism while preserving the asymptotic tail behavior of the observed data.
The choices of the censoring level and the contamination proportion are guided by both theoretical requirements and practical considerations. The proportion of uncensored extreme observations is fixed at $p\in\{0.55,0.7\},$ corresponding respectively to moderate and mild censoring regimes. Both values satisfy the weak censoring condition $p>1/2,$ which guarantees identifiability of the tail index and preserves regular variation of the observed data. Varying $p$ allows us to assess the impact of censoring severity on estimation accuracy while remaining within a statistically coherent framework.
@@ 4.3. Contamination schemes
To assess robustness, we consider contamination scenarios of increasing severity, ranging from the ideal uncontaminated setting to adverse configurations in which contamination affects the upper tail of the lifetime distribution. The contamination intensity is controlled by a parameter $\epsilon\in(0,1),$ which is common to all contaminated configurations.
In all contaminated settings, contamination is modeled through a mixture mechanism acting on the latent lifetime variable. Specifically, the distribution of $X$ is given by
$F_{\epsilon}=(1-\epsilon)F+\epsilon F_{c},$
where $F$ denotes the target distribution and $F_{c}$ is a heavier-tailed contaminating distribution with tail index $\gamma_{c}>\gamma_{1}.$ This mixture formulation offers a flexible and interpretable framework for robustness analysis, capturing the presence of extreme outliers while maintaining identifiability under independent censoring.
The condition $\gamma_{c}>\gamma_{1}$ ensures that contamination primarily affects the upper tail of the distribution, generating extreme observations that act as outliers relative to the target model. This setting reflects realistic departures from the ideal model and provides a stringent stress test for robustness. The resulting mixture formulation offers a flexible and interpretable framework for robustness analysis while preserving identifiability under independent censoring.
Several contamination configurations are considered:
In the uncontaminated case $(\epsilon=0),$ the lifetime variable $X$ and the censoring variable $C$ are generated independently from their target distributions $F$ and $G,$ respectively. The observed sample $(Z,\delta)$ is obtained through the standard random right-censoring mechanism. This configuration serves as a benchmark for assessing efficiency under correct model specification and for quantifying the efficiency loss induced by robustness-oriented procedures.
In the pre-censoring contamination setting, contamination is introduced at the level of the latent lifetime variable $X,$ prior to censoring. A proportion $\epsilon\in(0,1)$ of the $X$-values is replaced by extreme observations generated from a heavier-tailed distribution with tail index $\gamma_{c}>\gamma_{1}.$ The censoring variable $C$ is generated independently from its target distribution $G,$ whose tail behavior is governed by $\gamma_{2}.$ The censoring mechanism is then applied to the contaminated latent sample, yielding the observed pairs $(Z,\delta).$ This framework preserves the independent censoring structure and directly reflects the impact of contamination on the tail behavior of the lifetime distribution, making it a natural, coherent, and identifiable setting for robustness analysis in both parametric and nonparametric frameworks.
An alternative contamination mechanism consists in perturbing the observable pair $(Z,\delta)$ after the censoring mechanism has taken place. While such post-censoring contamination may be interpreted as data corruption or recording errors at the observational stage, it does not correspond to a well-defined censoring model in a nonparametric framework. By directly altering the joint distribution of $(Z,\delta),$ this mechanism breaks the link between the observed data and the underlying lifetime and censoring distributions $(F,G),$ thereby rendering reconstruction-based inference ill-posed. For this reason, post-censoring contamination is not retained in the simulation study and is viewed only as an illustrative stress mechanism rather than as a statistically coherent robustness framework.
Finally, contamination affecting the censoring variable $C$ is not considered. Such contamination would alter the censoring mechanism itself and may violate the independent censoring assumption underlying most theoretical results for censored extreme value estimation. In particular, contamination of $C$ directly modifies the joint distribution of the observable pair $(Z,\delta)$ in a way that cannot generally be associated with any well-defined latent lifetime distribution $F$. As a consequence, tail index estimation loses a clear statistical interpretation. Accordingly, robustness assessments in this paper focus exclusively on contamination acting on the lifetime variable prior to censoring.